OpenStudy (aykayyy):

how do i set this problem up? Use the method of cylindrical shells to find the volume gener- ated by rotating the region bounded by the given curves about the y-axis. Sketch the region and a typical shell. y=x^3, y=0, x=1; about y=1

OpenStudy (dumbcow):

 volume using shell method \[2\pi \int\limits radius*height\] because its being rotated around horizontal axis...radius must be in terms of "y" x = y^1/3 \[\rightarrow 2\pi \int\limits (1-y)(1-\sqrt[3]{y}) dy\]